solve the following application problem
the local movie theater sold 100 tickets for $390. If children's tickets cost $3 and adult tickets $6, how many of each were sold?
3c+6a=390
-3c -3c
6a=390-3c

Answer by Paul(988)

LEt the cjilderns tickets be x
let the adults tickets be y
Now x+y=100 Solve for y
y=100-x
Childern tickets = $3
Adults = $6
3x+6y=390 (SUBSITUTE)
3x+6(100-x)=390
3x+600-6x=390
-3x=-210
x=70
y=100-70
y=30

Hence, about 70 childern tickets were sold, and about 30 adults tickets were sold.
Paul.